# Enhanced Sampling for Efficient Learning of Coarse-Grained Machine Learning Potentials

### v1

@@ 1 Introduction
Traditional CG models have historically been parameterized using either ”top-down” or ”bottom-up” approaches. In top-down approaches, the model parameters are adjusted to reproduce macroscopic observables, such as experimental measurements. A well-known example is the MARTINI model, which was designed to reproduce experimental free energies 15. In bottom-up approaches 19, the goal is to construct a CG model that reproduces the thermodynamic/kinetic behavior of the fine-grained system. A common approach is to enforce thermodynamic consistency, ensuring that simulations with the CG model yield the same equilibrium distribution as fine-grained simulations projected onto the CG phase space 19, 12. By construction, the exact CG potential is the many-body potential of mean force (PMF). However, traditional attempts to approximate the PMF using functional forms or large basis sets similar to classical all-atom potentials have generally proven limited, as added complexity rarely guarantees improved accuracy or transferability 12.
@@
Another active line of research explores deep generative models for CG systems 38, 39, 40, 41, 42, 43. Boltzmann Emulators 44, 45, 46, for example, act as surrogate models by learning a biased distribution that enables one-shot sampling. The connection between generative models and molecular dynamics has led to new sampling approaches. For instance, Flow-Matching 47 improves data efficiency by training a normalizing flow to approximate the target distribution and then derives forces from the generated samples to train a CG MLP. This shares the goal of relative entropy minimization in reproducing the target distribution, but circumvents the need for iterative CG simulations. Diffusion models provide another approach by directly estimating forces via the score function to enable CG MD simulations 48, 49, 50, 51. Despite these advances, generative CG models face limitations: the lack of an explicit energy function prevents unbiased reweighting, scaling to larger systems remains difficult 52, 53, 54, 55, 56, 57, and training generally requires unbiased CG MD data. Energy-based models offer an alternative, since they do not rely on training samples 58, but typically require a reliable energy predictor 59, 60, 61, 62, 63, which in practice depends on the availability of existing CG MLPs 29.
@@ 2 Theory and Methods
The coarse-grained modeling begins with the definition of a mapping from the atomistic (AT) description to a reduced set of CG variables. Denote the AT coordinates as $\mathbf{r}\in\mathbb{R}^{3N}$ and the CG coordinates as $\mathbf{R}=\xi(\mathbf{r})\in\mathbb{R}^{3n}$, with $n<N$. The mapping $\xi$ groups atoms into beads, reducing dimensionality while providing a basis for constructing effective interactions that reproduce microscopic behavior. In this work, we focus on equilibrium thermodynamics in the canonical ($NVT$) ensemble and assume a linear and orthogonal mapping. Extensions to nonlinear mappings, non-equilibrium systems, and kinetic modeling have also been studied 64, 65, 66.
@@
A CG potential with parameters $\theta$, denoted $U(\mathbf{R};\theta)$, is thermodynamically consistent if
$U(\mathbf{R};\theta)=-k_{B}T\ln p_{\text{CG}}(\mathbf{R})+\text{constant},$ (3)
Since the partition function $\mathcal{Z}$ is generally intractable, schemes such as force matching (FM) or relative entropy minimization are typically used to learn $U(\mathbf{R};\theta)$.
@@ 2.1 Force Matching
$\hat{\chi}^{2}(\theta)=\frac{1}{3nM}\sum_{i=1}^{M}\Big\|\,\xi(\mathbf{f}(\mathbf{r}_{i}))+\nabla U(\xi(\mathbf{r}_{i});\theta)\,\Big\|^{2},$ (6)
@@ 2.2 Finite Data Size Effects
The first arises from the nature of CG force matching. Unlike atomistic MLPs, where potential energy labels are directly available, the CG PMF $U(\mathbf{R})$ must be inferred indirectly from instantaneous forces by minimizing the variational bound in Eq. 4. The true mean force $\mathbf{F}(\mathbf{R})$ is defined as the average of all atomistic configurations corresponding to the same coarse-grained state $\mathbf{R}$, while a single projected force represents only one noisy sample of this average. Accurate approximation of this mean requires dense sampling in the neighborhood of each $\mathbf{R}$, so that statistical noise does not dominate the learning signal. As a result, FM-trained CG models generally require much larger datasets than atomistic models trained on explicit energy surfaces. Although conditional averages could, in principle, be obtained using constrained MD or Blue Moon sampling 67, 68, performing such targeted sampling for a sufficiently dense set of coarse-grained configurations is computationally prohibitive.
@@ 2.3 Unbiased Mean Forces from Biased Sampling
where $\beta=(k_{B}T)^{-1}$ and $\mathcal{Z}_{W}$ is the biased partition function. The conditional distribution of atomistic configurations given a CG configuration $\mathbf{R}$ is unaffected by a bias that depends only on $\mathbf{R}$. This follows because the bias factor cancels when conditioning on $\mathbf{R}$, leaving the conditional distribution identical to that of the unbiased ensemble (see SI for a detailed derivation). Consequently, the conditional expectation of the projected forces is also unchanged:
@@
Recomputing biased instantaneous forces with respect to the unbiased atomistic potential $u(\mathbf{r})$ therefore yields the correct unbiased mean forces. Intuitively, the bias alters how frequently a given CG configuration $\mathbf{R}$ is visited, but once $\mathbf{R}$ is fixed, the conditional distribution of atomistic microstates consistent with $\mathbf{R}$ is unaffected by $W$.
Equation 8 shows that the mean force $\mathbf{F}(\mathbf{R})$ is invariant under CG coordinate-dependent biasing, provided that forces are evaluated from the unbiased atomistic potential. Consequently, the FM objective can be expressed as an expectation over the biased distribution:
$\chi^{2}(\theta)=\Big\langle\big\|\,\xi(\mathbf{f}(\mathbf{r}))+\nabla U(\xi(\mathbf{r});\theta)\,\big\|^{2}\Big\rangle_{\mathbf{r_{W}}}.$ (9)
This invariance enables the training of CG potentials on datasets generated with biased sampling, without requiring reweighting of the loss function. The practical advantages are twofold: (i) biased simulations accelerate exploration of rarely visited states, reducing the total simulation time needed for data generation, and (ii) the resulting datasets provide more uniform coverage of both energy basins and transition regions, leading to more accurate and robust CG MLPs. An overview of the enhanced sampling force matching workflow is shown in Figure 1.
Figure 1: Overview of the enhanced sampling force matching method. (A) Classical force matching: positions $\mathbf{r}$ and forces $\mathbf{f}$ from an unbiased atomistic MD simulation are used to learn the potential of mean force (PMF) $U$. (B) Enhanced sampling force matching (this work): Configurations are obtained via enhanced sampling, reducing the required simulation time (light red region). Unbiased forces, $\mathbf{f}_{W}$, are then recomputed using the unbiased potential, which incurs minimal additional computational cost. The PMF is learned from the biased configurations, $\mathbf{r}_{W}$, and their corresponding recomputed forces, $\mathbf{f}_{W}$.
@@ 2.4 Enhanced Sampling Methods
In plain metadynamics, the bias keeps growing indefinitely, eventually flattening the free energy surface. Well-Tempered metadynamics 69 improves this by tempering the Gaussian heights with a bias factor $\gamma>1$,
@@
where $A(\mathbf{R})$ is the free energy surface of the CG (or CV) coordinates R. This corresponds to sampling at an effective temperature $T^{*}=\gamma T$, since the bias factor $\gamma=T^{*}/T$ rescales the thermal fluctuations along $\mathbf{R}$. The term “well-tempered” reflects the fact that the bias is added more slowly over time, striking a balance between the exploration of new regions and the preservation of meaningful free energy differences.
@@ 2.5 Graph Neural Network Potentials without Priors
To parameterize the CG potential $U(\mathbf{R};\theta)$ for molecular systems, we adopt the MACE architecture 70, an equivariant message-passing graph neural network originally developed for atomistic potentials. Each CG bead is represented as a node in a graph, and edges indicate neighbor pairs within a cutoff radius and carry distance/relative-vector embeddings. Equivariant message passing layers update node features while enforcing that $U(\mathbf{R};\theta)$ is invariant under rigid translations and rotations and that internal vector features transform equivariantly. The force on bead $i$ is obtained directly from the learned potential by automatic differentiation,
@@
A common strategy in the literature is to augment CG MLPs with a physics-based prior, such as baseline pairwise interactions and harmonic bonded terms, so that the network only learns a corrective energy term, which is helpful for data efficiency and improves simulation stability. In contrast, here we train MACE directly on the force-matching loss (Eq. 4) without including any prior 71. This choice avoids introducing modeling bias and allows the network to learn the PMF purely from data.
@@ 3.1 Low-dimensional Müller Brown Potential
We consider the two-dimensional Müller–Brown (MB) potential, a canonical test system for transition path sampling 72, which features a global minimum and two local minima separated by saddle points (Figure 2A).
@@
Figure 4A reports the mean-squared error (MSE) between the predicted and exact mean force as a function of the amount of training data. Models trained on biased datasets reach lower error and variance with only a few thousand samples, whereas models trained on unbiased datasets require orders of magnitude more data, yet still exhibit larger variance and higher error. Direct comparison of the learned force curves (Figure 4B–C) further illustrates this difference: While both models reproduce the mean force in densely sampled regions, biased training achieves much lower uncertainty and accurately recovers both the overall shape and the fine features of the mean force with limited data. Unbiased training, on the contrary, captures only the broad trend and fails to reproduce local structure even with orders of magnitude more samples.
@@
We then train the MACE model on these datasets using the chemtrain framework 75, 76, with the same hyperparameter settings (listed in the SI). CG simulations are performed under Langevin dynamics at 300 K using JAX M.D. 77. For evaluation, we run $100$ independent CG simulations of $100$ ns each, initialized from random configurations, for both the unbiased dataset and biased datasets with different bias factors $\gamma$. Ramachandran plots and Dihedral distributions (Figure 5B) show that models trained on unbiased data fail to recover the metastable basin $\alpha_{\mathrm{L}}$ at $\phi\approx 0^{\circ}$$100^{\circ}$ on the right-hand side of the Ramachandran map, whereas biased training with sufficiently large $\gamma$ recovers both modes accurately. Quantifying metastable populations across five independent models (Supporting Information Figure S6-7) show that unbiased datasets and low-$\gamma$ WT MetaD assign nearly zero probability to the metastable state $\alpha_{\mathrm{L}}$, while higher-$\gamma$ datasets accurately capture it.
@@
Finally, we assess stability by monitoring numerical instabilities across training fractions (Figure 6C). Chains are considered unstable and removed when the predicted potential energy reaches unphysically high values. In previous analyses, these unstable chains were already excluded; here we explicitly report their occurrence. Most simulations remain stable across 100 ns, with failures occurring only in a few chains. For the 500 ns unbiased dataset, up to six chains diverge at fraction 0.75, with isolated failures at fractions 0.2, 0.3, and 1.0. The 2 µs unbiased dataset shows single-chain failures at fractions 0.4 and 0.75. In contrast, biased datasets exhibit only one failure at fraction 0.3. These results indicate that biased datasets improve both accuracy and stability by providing broader coverage of configuration space. For completeness, we report free energy surfaces without chain removal as well as per-chain results (SI, Figure S8-9).

### v2

@@ 1 Introduction
Traditional CG models have historically been parameterized using either ”top-down” or ”bottom-up” approaches. In top-down approaches, the model parameters are adjusted to reproduce macroscopic observables, such as experimental measurements. A well-known example is the MARTINI model, which was designed to reproduce experimental partition coefficients 15. In bottom-up approaches 19, the primary goal is to construct a CG model that reproduces the equilibrium configurational distribution (or free energy landscape) of the fine-grained system 12. While this ensures the model captures correct static correlations, it does not guarantee the accurate reproduction of dynamical properties. 19, 12. By construction, the exact CG potential is the many-body potential of mean force (PMF). However, traditional attempts to approximate the PMF using functional forms or large basis sets similar to classical all-atom potentials have generally proven limited, as added complexity rarely guarantees improved accuracy or transferability 12.
@@
Another active line of research explores deep generative models for CG systems 38, 39, 40, 41, 42, 43, 44. Boltzmann Emulators 45, 46, 47, for example, act as surrogate models by learning a biased distribution that enables one-shot sampling. The connection between generative models and molecular dynamics has led to new sampling approaches. For instance, Flow-Matching 48 improves data efficiency by training a normalizing flow to approximate the target distribution and then derives forces from the generated samples to train a CG MLP. This shares the goal of relative entropy minimization in reproducing the target distribution, but circumvents the need for iterative CG simulations. Diffusion models provide another approach by directly estimating forces via the score function to enable CG MD simulations 49, 50, 51, 52, 53. Despite these advances, generative CG models face limitations: the lack of an explicit energy function prevents unbiased reweighting, scaling to larger systems remains difficult 54, 55, 56, 57, 58, 59, and training generally requires unbiased CG MD data. Energy-based models offer an alternative 60, since they do not rely on training samples, but typically require a reliable energy predictor 61, 62, 63, 64, 65, which in practice depends on the availability of existing CG MLPs 29.
@@ 2 Theory and Methods
The coarse-grained modeling begins with the definition of a mapping from the atomistic (AT) description to a reduced set of CG variables. Denote the AT coordinates as $\mathbf{r}\in\mathbb{R}^{3n}$ and the CG coordinates as $\mathbf{R}=\xi(\mathbf{r})\in\mathbb{R}^{3N}$, with $N<n$. The mapping $\xi$ groups atoms into beads, reducing dimensionality while providing a basis for constructing effective interactions that reproduce microscopic behavior. In this work, we focus on equilibrium thermodynamics in the canonical ($NVT$) ensemble and assume a linear and orthogonal mapping. Extensions to nonlinear mappings, non-equilibrium systems, and kinetic modeling have also been studied 66, 67, 68.
@@
The exact many-body potential of mean force (PMF) $U^{*}(\mathbf{R})$ is defined by the relation:
$U^{*}(\mathbf{R})=-k_{B}T\ln p_{\text{CG}}(\mathbf{R})+C,$ (3)
where $C$ is an arbitrary additive constant. Since exact marginalization is generally intractable for complex systems, a parametric model $U(\mathbf{R};\theta)$ is typically learned to approximate the exact PMF. The parameters $\theta$ are determined by minimizing a variational objective, such as force matching or relative entropy minimization, such that $U(\mathbf{R};\theta)\approx U^{*}(\mathbf{R})$.
@@ 2.1 Force Matching
$\hat{\chi}^{2}(\theta)=\frac{1}{3M}\sum_{i=1}^{M}\Big\|\,\xi(\mathbf{f}(\mathbf{r}_{i}))+\nabla U(\xi(\mathbf{r}_{i});\theta)\,\Big\|^{2},$ (6)
@@ 2.2 Finite Data Size Effects
The first arises from the nature of CG force matching. Unlike atomistic MLPs, where potential energy labels are directly available, the CG PMF $U(\mathbf{R})$ must be inferred indirectly from instantaneous forces by minimizing the variational bound in Eq. 4. The true mean force $\mathbf{F}(\mathbf{R})$ is defined as the average of all atomistic configurations corresponding to the same coarse-grained state $\mathbf{R}$, while a single projected force represents only one noisy sample of this average. Accurate approximation of this mean requires dense sampling in the neighborhood of each $\mathbf{R}$, so that statistical noise does not dominate the learning signal. As a result, FM-trained CG models generally require much larger datasets than atomistic models trained on explicit energy surfaces. Although conditional averages could, in principle, be obtained using constrained MD or Blue Moon sampling 69, 70, performing such targeted sampling for a sufficiently dense set of coarse-grained configurations is computationally prohibitive.
@@ 2.3 Unbiased Mean Forces from Biased Sampling
The key observation is that although the bias changes the marginal distribution of the CG coordinates, it does not alter the conditional distribution of atomistic configurations at fixed $\mathbf{R}$. The conditional distribution under the bias is
$p_{W}(\mathbf{r}\mid\mathbf{R})\;=\;\frac{\delta(\xi(\mathbf{r})-\mathbf{R})\,e^{-\beta\big(u(\mathbf{r})+W(\mathbf{R})\big)}}{\int\delta(\xi(\mathbf{r})-\mathbf{R})\,e^{-\beta\big(u(\mathbf{r})+W(\mathbf{R})\big)}\,d\mathbf{r}}.$ (8)
Since $W(\mathbf{R})$ is constant when conditioning on $\mathbf{R}$, it cancels between numerator and denominator:
$p_{W}(\mathbf{r}\mid\mathbf{R})\;=\;\frac{\delta(\xi(\mathbf{r})-\mathbf{R})\,e^{-\beta u(\mathbf{r})}}{\int\delta(\xi(\mathbf{r})-\mathbf{R})\,e^{-\beta u(\mathbf{r})}\,d\mathbf{r}}\;=\;p(\mathbf{r}\mid\mathbf{R}),$ (9)
Thus, any bias that depends exclusively on the CG variables leaves the atomistic conditional ensemble at fixed $\mathbf{R}$ unchanged. Consequently, the coarse-grained mean force, defined as the conditional expectation of the projected forces, remains invariant:
@@
Intuitively, the bias alters how frequently a given CG configuration $\mathbf{R}$ is visited, but once $\mathbf{R}$ is fixed, the conditional distribution of atomistic microstates consistent with $\mathbf{R}$ is unaffected by $W$. In practice, however, the invariance derived in Eq. 10 relies on several underlying assumptions, outlined below:
  1. 1.
    Functional Dependence of the Bias: The bias potential must depend exclusively on the CG variables, i.e., $W=W(\xi(\mathbf{r}))$. Dependence on orthogonal degrees of freedom would perturb the conditional ensemble $p(\mathbf{r}|\mathbf{R})$, rendering the sampled mean forces thermodynamically inconsistent. In such cases, explicit reweighting is required to recover unbiased mean forces.
  2. 2.
    Conditional Ergodicity: The system needs to remain ergodic in the orthogonal degrees of freedom. While the bias accelerates the exploration of $\mathbf{R}$, the fast degrees of freedom (e.g., vibrations, solvent) are assumed to equilibrate rapidly to the conditional Boltzmann distribution. Hysteresis or trapping in these hidden variables would yield non-equilibrium force estimates.
  3. 3.
    Recovery of Unbiased Forces: The training targets must be the forces derived from the unbiased potential $u(\mathbf{r})$. The biasing force contribution, $\nabla_{\mathbf{r}}W(\xi(\mathbf{r}))$, must be subtracted from biased forces.
Subject to the assumptions above, the optimization over the biased distribution targets the identical mean force as the unbiased case. The biased objective function is defined as:
$\chi^{2}_{W}(\theta)\;=\;\Big\langle\big\|\,\xi(\mathbf{f}(\mathbf{r}))+\nabla U(\xi(\mathbf{r});\theta)\,\big\|^{2}\Big\rangle_{\mathbf{r_{W}}}.$ (11)
While the scalar value of the loss $\chi^{2}_{W}$ differs from the unbiased objective $\chi^{2}$ due to the modified sampling density, both functionals share the same global minimum with respect to $\theta$ (see SI for a detailed proof). This invariance enables the training of CG potentials on datasets generated with biased sampling, without requiring reweighting of the loss function. The practical advantages are twofold: (i) biased simulations accelerate exploration of rarely visited states, reducing the total simulation time needed for data generation, and (ii) the resulting datasets provide more uniform coverage of both energy basins and transition regions, leading to more accurate and robust CG MLPs. An overview of the enhanced sampling force matching workflow is shown in Figure 1.
Figure 1: Overview of the enhanced sampling force matching method. (A) Classical force matching: positions $\mathbf{r}$ and forces $\mathbf{f}$ from an unbiased atomistic MD simulation are used to learn the potential of mean force (PMF) $U$. (B) Enhanced sampling force matching (this work): Configurations are obtained via enhanced sampling, reducing the required simulation time (light red region). The forces acting during the biased simulation are $\hat{\mathbf{f}}_{W}(\mathbf{r}_{W})=-\nabla_{\mathbf{r}}\!\left(u(\mathbf{r}_{W})+W(\xi(\mathbf{r}_{W}))\right)$, while the unbiased forces used for training are recomputed using the unbiased potential, $\mathbf{f}_{W}(\mathbf{r}_{W})=-\nabla_{\mathbf{r}}u(\mathbf{r}_{W})$, which incurs minimal additional computational cost. The PMF is learned from the biased configurations $\mathbf{r}_{W}$ and their corresponding recomputed forces $\mathbf{f}_{W}$.
@@ 2.4 Enhanced Sampling Methods
In plain metadynamics, the bias keeps growing indefinitely, eventually flattening the free energy surface. Well-Tempered metadynamics 71 improves this by tempering the Gaussian heights with a bias factor $\gamma>1$,
@@
where $A(\mathbf{R})$ is the free energy surface of the CG (or CV) coordinates R. This corresponds to sampling at an effective temperature $T^{*}=\gamma T$, since the bias factor $\gamma=T^{*}/T$ rescales the thermal fluctuations along $\mathbf{R}$. Physically, this implies that the coarse-grained variables explore the landscape as if coupled to a high-temperature heat bath at $T^{*}$, effectively reducing barrier heights by a factor of $1/\gamma$. Note that the remaining atomistic degrees of freedom (e.g., solvent, fast vibrations) remain at the physical temperature $T$. This allows the system to overcome high-energy barriers associated with the collective variables without inducing the unphysical structural denaturation that would occur in a global high-temperature simulation. The term “well-tempered” reflects the fact that the bias is added more slowly over time, striking a balance between the exploration of new regions and the preservation of meaningful free energy differences.
A practical challenge in applying enhanced sampling to novel systems is selecting bias parameters without prior knowledge of the free-energy landscape. Importantly, our method prioritizes achieving sufficient exploration of the conformational space rather than the exact convergence of the bias potential required for quantitative free-energy reconstruction. This distinction makes the method more robust to sub-optimal parameter choices. For novel systems, we recommend an iterative refinement strategy. Initial parameters can be selected based on standard heuristics 71, 72, 73: for solvated biomolecules, the bias factor is typically chosen in the range $\gamma\in[5,20]$ to adequately reduce expected barriers, while the Gaussian width $\sigma$ is estimated from natural thermal fluctuations observed in short unbiased precursor runs (e.g., $\sigma\approx 1/3$74. Sampling quality is then assessed by monitoring the histogram of the collective variables. If unsampled regions or “gaps” persist along transition pathways, the bias strength can be increased to help bridge these regions.
@@ 2.5 Graph Neural Network Potentials without Priors
To parameterize the CG potential $U(\mathbf{R};\theta)$ for molecular systems, we adopt the MACE architecture 75, an equivariant message-passing graph neural network originally developed for atomistic potentials. Each CG bead is represented as a node in a graph, and edges indicate neighbor pairs within a cutoff radius and carry distance/relative-vector embeddings. Equivariant message passing layers update node features while enforcing that $U(\mathbf{R};\theta)$ is invariant under rigid translations and rotations and that internal vector features transform equivariantly. The force on bead $i$ is obtained directly from the learned potential by automatic differentiation,
@@
A common strategy in the literature is to augment CG MLPs with a physics-based prior, such as baseline pairwise interactions and harmonic bonded terms, so that the network only learns a corrective energy term, which is helpful for data efficiency and improves simulation stability. In contrast, here we train MACE directly on the force-matching loss (Eq. 4) without including any prior 76. This choice avoids introducing modeling bias and allows the network to learn the PMF purely from data.
@@ 3.1 Low-dimensional Müller Brown Potential
We consider the two-dimensional Müller–Brown (MB) potential, a canonical test system for transition path sampling 77, which features a global minimum and two local minima separated by saddle points (Figure 2A).
@@
Figure 4A reports the mean-squared error (MSE) between the predicted and exact mean force as a function of the amount of training data. Models trained on biased datasets reach lower error and variance with only a few thousand samples, whereas models trained on unbiased datasets require orders of magnitude more data, yet still exhibit larger variance and higher error. Direct comparison of the learned force curves (Figure 4B–C) further illustrates this difference: While both models reproduce the mean force in densely sampled regions, biased training achieves much lower uncertainty and accurately recovers both the overall shape and the fine features of the mean force with limited data. Unbiased training, on the contrary, captures only the broad trend and fails to reproduce local structure even with orders of magnitude more samples. We additionally compare the exact free energy profile with MLP simulations in the Supporting Information (Figure S2), where simulation profiles are obtained via the negative natural logarithm of the sampled position histograms.
@@
We then train the MACE model on these datasets using the chemtrain framework 79, 80, with the same hyperparameter settings (listed in the SI). CG simulations are performed under Langevin dynamics at 300 K using JAX M.D. 81. For evaluation, we run $100$ independent CG simulations of $100$ ns each, initialized from random configurations, for both the unbiased dataset and biased datasets with different bias factors $\gamma$. Ramachandran plots and Dihedral distributions (Figure 5B) show that models trained on unbiased data fail to recover the metastable basin $\alpha_{\mathrm{L}}$ at $\phi\approx 0^{\circ}$$100^{\circ}$ on the right-hand side of the Ramachandran map, whereas biased training with sufficiently large $\gamma$ recovers both modes accurately. Quantifying metastable populations across five independent models (Supporting Information Figure S7-8) show that unbiased datasets and low-$\gamma$ WT MetaD assign nearly zero probability to the metastable state $\alpha_{\mathrm{L}}$, while higher-$\gamma$ datasets accurately capture it.
@@
Finally, we assess stability by monitoring numerical instabilities across training fractions (Figure 6C). Chains are considered unstable and removed when the predicted potential energy reaches unphysically high values. In previous analyses, these unstable chains were already excluded; here we explicitly report their occurrence. Most simulations remain stable across 100 ns, with failures occurring only in a few chains. For the 500 ns unbiased dataset, up to six chains diverge at fraction 0.75, with isolated failures at fractions 0.2, 0.3, and 1.0. The 2 µs unbiased dataset shows single-chain failures at fractions 0.4 and 0.75. In contrast, biased datasets exhibit only one failure at fraction 0.3. These results indicate that biased datasets improve both accuracy and stability by providing broader coverage of configuration space. For completeness, we report free energy surfaces without chain removal as well as per-chain results (SI, Figure S9-10). Additionally, to assess computational efficiency, we present a wall-clock time comparison between our MACE model and the reference ATMD simulations performed in GROMACS (SI, Table A2).